OpenStudy (anonymous):

If f(x) = x 2 + 1 and g(x) = 3x + 1, find g(20) + f(6).

OpenStudy (anonymous):

f(x)= x^2+1 f(6)=(6)^2 +1 =37 g(x)=3x+1 g(20)= 3(20)+1 =61 g(20)+ f(6) = 61+37 =98 but i'm not sure whether f(x)= x^2 +1